# Think Smart, Not Hard: Difficulty Adaptive Reasoning for Large Audio Language Models

### v1

@@ 1 Introduction
In recent years, general artificial intelligence advances rapidly with the development of large language models (LLMs) (Dubey et al., 2024; Hurst et al., 2024; Team et al., 2024; Team, 2024). The reasoning ability of LLMs is further enhanced by the chain-of-thought (CoT) paradigm (Jaech et al., 2024; DeepSeek-AI et al., 2025; Yang et al., 2025), which significantly improves performance on complex problems. At the same time, Large Audio Language Models (LALMs) (Tang et al., 2024; Chu et al., 2023; 2024) also progress rapidly and raise an important question: how can they achieve reasoning more efficiently and effectively?
@@
In previous work, Xie et al. (2025) enable reasoning in LALMs through supervised fine-tuning (SFT) on a large-scale dataset with CoT annotations, providing the first evidence that CoT is effective for solving complex audio understanding problems. Building on this, Li et al. (2025) introduces group relative policy optimization (GRPO) into LALMs for the first time, surpassing the former with less data and offering initial comparisons across different prompts. These studies focus either on SFT or GRPO, and both apply reasoning to all questions, but they lack in-depth analysis of the differences between the two approaches. Therefore, we conduct a detailed analysis to determine under what conditions SFT and GRPO are more effective. We find that GRPO performs better on harder questions, whereas on easier ones it tends to produce redundant reasoning and performs slightly worse than SFT. Furthermore, we analyze GRPO under forced reasoning (explicit prompt) and without forced reasoning (implicit prompt), and observe that the forced reasoning models maintain a clear advantage on harder questions, while the two settings perform similarly on easier ones. Taken together, these findings indicate that achieving efficient reasoning in LALMs requires adapting reasoning length to problem difficulty.
@@
Regarding reasoning length and efficiency, Qu et al. (2025) highlight issues such as redundancy and overthinking. RL-based studies on LLMs (Arora & Zanette, 2025; Aggarwal & Welleck, 2025) design length-penalty rewards but rely on fixed thresholds that overlook question types and difficulty levels. In the audio domain, Wu et al. (2025) introduces a “when to think” mechanism that guides the model on whether reasoning is necessary. However, for the samples that still perform reasoning, it lacks a fine-grained mechanism to modulate how much to reason.Therefore, an effective and efficient LALM should reason smartly by adapting its reasoning depth to the problem’s complexity, achieving short reasoning for simple questions and deeper reasoning for difficult ones. Based on this analysis and prior work, we propose a new length-based reward function that no longer depends on fixed thresholds. In addition, we introduce two difficulty-adaptive standards to complement this reward, enabling reasoning length to align more appropriately with question difficulty. Both approaches achieve strong performance on the MMAU benchmark (Sakshi et al., 2025), particularly on harder questions, while also producing much shorter reasoning than direct GRPO models, thereby greatly improving reasoning efficiency. Furthermore, we conduct a qualitative case study on models from the main experiments, providing a detailed analysis of the reasoning structures in their outputs.
@@ 2 Deep Analysis of different methods for LALMs
LALMs have already demonstrated strong capability in addressing basic understanding tasks. Recent studies (Xie et al., 2025; Li et al., 2025) focus on enhancing their ability to solve complex problems through reasoning. However, while these works adopt different implementation methods, they lack in-depth analysis of the differences between approaches. Therefore, in this section, we conduct detailed analysis from two perspectives: which is more effective under various conditions between SFT and GRPO, and whether performance gains are driven by explicit or implicit prompt.
@@
Audio Grounding AudioGrounding (Xu et al., 2021) 1805
Sound Classification VocalSound (Gong et al., 2022) 15531
TUT2017 (Mesaros et al., 2017) 3744
Sound Question Answering Clotho-AQA (Lipping et al., 2022) 6615
AVQA (Yang et al., 2022) 36036
@@
For the base models, we primarily adopt Qwen2-Audio-7B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2-Audio-7B-Instruct and Qwen2.5-Omni-7B  [^2] https://huggingface.co/Qwen/Qwen2.5-Omni-7B. More detailed experimental settings, including the two prompts, are provided in Appendix  A.1. At the same time, for those experiments that share common settings, we also add an additional group average in Table  2 to facilitate comparisons across groups.
@@ 3.1 Defining Model-perspective Difficulty
$C=\sum_{i=1}^{G}c_{i}$ (1)
$\gamma=\begin{cases}0,&C\geq 6,\\ 0.5,&3\leq C<6,\\ 1,&C<3,\end{cases}\ (G=8)$ (2)
@@
$a^{(n)}_{j}=\mathbf{A}^{(n)}_{T,j},\ j=1,\cdots,T$ (3)
$\bar{p}_{j}=\frac{1}{N}\sum_{n=1}^{N}a^{(n)}_{j}\ ,\ H=-\sum\bar{p}_{j}\log\bar{p}_{j}\ ,(j\in\mathcal{M})$ (4)
$\gamma^{(b)}=\frac{H^{(b)}-\min_{b^{\prime}\in\mathcal{B}}H^{(b^{\prime})}}{\max_{b^{\prime}\in\mathcal{B}}H^{(b^{\prime})}-\min_{b^{\prime}\in\mathcal{B}}H^{(b^{\prime})}}\;\in\;[0,1]$ (5)
Here, $\mathbf{A}$ denotes the complete attention matrix, $T$ represents the number of tokens, $\mathcal{M}$ indicates the indices corresponding to the audio attention part, $H$ denotes the computed entropy, and $\mathcal{B}$ is the batch size. The final $\gamma^{(b)}$ represents the difficulty value of the $b$-th sample within a batch. Compared with the first method, this approach provides a more fine-grained difficulty division, no longer limited to three categories, but instead yielding a continuous difficulty range.
@@ 3.2 Difficulty-Adaptive Length-Based Reward
After defining question difficulty, the next step is to link difficulty with reasoning length in an effective way. Our core idea is short reasoning for simple questions and long reasoning for difficult ones, with reward values that change dynamically rather than staying fixed. Thus, we design a rule-based reward using a negative exponential function, with the corresponding curve shown in Figure  1. Specifically, it smoothly adjusts reward values according to the ratio between the current reasoning length and the model’s maximum reasoning length. The decision between reward and penalty also depends on correctness. Incorrect samples always receive a penalty that decreases as reasoning length grows, encouraging further reasoning. Once a sample is answered correctly at a certain stage, the penalty switches to a positive reward, with the reward value increasing as reasoning length shortens, encouraging concise reasoning that retains only the core solution steps. For difficulty, simple questions correspond to steeper curves, while difficult questions correspond to flatter curves, meaning that for the same reward value, difficult questions correspond to longer reasoning lengths, and for the same reasoning length, simple questions yield smaller rewards. This achieves short reasoning for simple tasks and longer reasoning for difficult ones. The detailed calculation is as follows:
@@
Figure 1: Curves of GRDR and GA2DR with normalized length.
@@ 4 Experiments
Setup.
In this section, our experiments are mainly based on Qwen2-Audio-7B-Instruct and Qwen2.5-Omni-7B. All models are fine-tuned using LoRA with Prompt2. All evaluations are conducted on MMAU-test-mini with ACC as the metric. Other details can be found in Appendix  A.1.
Table 3: The performance of models on MMAU-test-mini with different base models and reward settings, reported under both the human-perspective difficulty annotations and the assigned model-perspective difficulty annotations. Here, $\dagger$ denotes the model-perspective difficulty annotations.
@@
Gemini2.5-Pro-0506111https://ai.google.dev/gemini-api/docs/models#gemini-2.5-pro 70.57 65.26 62.16 52.32 95.82 77.64 57.47 55.60 12.35 66.00
@@ 4.1 Main Results
Table  3 mainly presents three parts: the performance of four baseline models and our proposed rewards on Qwen2-Audio-7B-Instruct and Qwen2.5-Omni-7B. The first column lists the different models. TR refers to the basic Truncation Reward (Liu et al., 2025), with its implementation formula provided in Appendix  A.1. In the experiments, truncation lengths are set to 120 and 400 tokens for Qwen2-Audio and Qwen2.5-Omni, respectively, based on the maximum length of correct samples from direct GRPO. Lengths are measured in tokens using the Qwen2-Audio tokenizer. In addition, we supplement MMAU-test-mini with model-perspective difficulty labels annotated by the four baseline models shown in Table  3, with the details described in Appendix  A.2.
@@
First, we analyze the four “Baseline Models”. Their performance improves from top to bottom, reflecting different capability levels and the average performance of most models, which makes them suitable for defining model-perspective difficulty labels. All models show consistent trends on easy and medium questions, but their performance diverges significantly on hard ones. We argue that easy and medium questions rely more on the LLM backbone within the LALM, as seen when comparing “Qwen2-Audio-7B-Instruct” with “Qwen2.5-Omni-7B”. In contrast, hard questions demand stronger audio information extraction, giving the former an advantage as a model dedicated solely to the audio modality.
Second, we compare our proposed GRDR and GA2DR with TR. In the “Based on Qwen2-Audio-7B-Instruct” part, overall performance across the three models is similar, with GRDR performing best and TR worst, proving the effectiveness of our method. In detail, TR performs better than GRDR on medium questions, but GRDR is stronger on hard ones. In the “Based on Qwen2.5-Omni-7B” part, both of our methods show clear overall improvements. More specifically, they perform much better than TR on medium and hard questions, while performance on easy questions is close. This demonstrates that our methods can effectively leverage question difficulty to assign suitable rewards, encourage deeper reasoning for hard questions, and better handle difficult tasks. Unlike TR, which relies on a fixed length threshold that only constrains part of the samples, our methods can balance questions across different difficulty levels and achieve stronger overall performance.
Third, we compare GRDR and GA2DR, focusing on the “Based on Qwen2.5-Omni-7B” part. GRDR performs better on easy questions, GA2DR is better on medium ones, and the two are close on hard questions. We believe this is because of their different difficulty standards: GRDR uses only three levels, while GA2DR does not limit difficulty grading and can more finely separate question difficulty, which is especially helpful for medium questions. For easy and hard questions, GA2DR is weaker mainly because some samples with similar audio attention entropy are split into different difficulty levels after normalization, reducing its advantage.
@@ 4.2 Analysis of Reasoning Length across Models
In this subsection, we mainly compare reasoning length under both perspectives of difficulty, as shown in Figure  2, focusing on the four baseline models and all models based on Qwen2.5-Omni-7B. Since Gemini2.5-Pro-0506 produces much longer reasoning than all other models, the lengths in the figure are shown after logarithmic transformation.
@@
4.3 Case Study: Qualitative Analysis of Output Paradigms across Models
In this section, we analyze output paradigms across models using 20 MMAU-test-mini questions on Dissonant Emotion Interpretation, focusing on sarcasm cause detection. The corresponding audios often involve multiple speakers and complex environments, making the task challenging and well-suited for qualitative analysis of different model outputs. In this section, in addition to the main models discussed above, we also include a Cold-Start GRPO model based on Qwen2-Audio-7B-Instruct, with detailed settings provided in the Appendix  A.1.
The detailed results are shown in Table  4, which include ACC and reasoning length on the subset mentioned above. In the second part, “Based on Qwen2-Audio-7B-Instruct,” we compare our three GRPO models with Audio-Reasoner. GRPO and GRPO with GRDR match Audio-Reasoner in performance while producing much shorter outputs, whereas the GRPO model after Cold-Start produces reasoning lengths close to Audio-Reasoner but performs worst. We believe this is mainly because the base model is weaker and can only imitate the surface structure of advanced model paradigms, showing that learning paradigms through Cold-Start is not effective in all cases and can even yield negative effects. In the “Based on Qwen2.5-Omni-7B” part of the table, performance differences are minor, mainly because the Qwen2.5-Omni base model already has some reasoning ability and shows good performance on complex questions. However, with our two proposed rewards, the models maintain their original performance while significantly shortening reasoning length, greatly improving reasoning efficiency.
Model ACC Avg-Length
Advanced Proprietary Model
Gemini2.5-Pro-0506 95 931.8
Based On Qwen2-Audio-7B-Instruct
Audio-Reasoner 75 547.1
GRPO 75 50.7
  $x$ + GRDR 70 38.5
  $x$ + Cold-Start SFT 55 541.4
Based On Qwen2.5-Omni-7B
GRPO 100 109.5
  $x$ + GRDR 95 56.2
  $x$ + GRDR ($l_{min}=0.1$) 90 94.1
  $x$ + GA2DR 100 56.9
  $x$ + GA2DR ($l_{min}=0.1$) 95 94.2
In addition to analyzing performance and reasoning length above, we also compared the outputs of different models on these 20 questions in detail. From the results, we found that models with better performance usually follow a complete reasoning process, which includes first grounding the audio and providing a corresponding caption—equivalent to identifying the known conditions in the problem—then reasoning based on these conditions, analyzing each option step by step, and finally giving the answer. This approach helps LALMs make better use of known information together with their broad pretrained knowledge to perform reasonable reasoning while also improving readability. However, weaker models, even when using this structure, often make mistakes in the first step of extracting known information. These errors propagate through the reasoning process, leading to wrong answers and lower performance. For this part, we also provide an example of different models’ outputs in Appendix  A.5, Figure  5 as reference.

### v2

@@ 1 Introduction
In recent years, general artificial intelligence advances rapidly with the development of large language models (LLMs) (dubey2024llama; hurst2024gpt; team2024gemini; team2024qwen2). The reasoning ability of LLMs is further enhanced by the chain-of-thought (CoT) paradigm (DBLP:journals/corr/abs-2412-16720; DBLP:journals/corr/abs-2501-12948; DBLP:journals/corr/abs-2505-09388), which significantly improves performance on complex problems. At the same time, Large Audio Language Models (LALMs) (DBLP:conf/iclr/TangYSC [^0] M024; DBLP:journals/corr/abs-2311-07919; DBLP:journals/corr/abs-2407-10759) also progress rapidly and raise an important question: how can they achieve reasoning more efficiently and effectively?
@@
In previous work, DBLP:journals/corr/abs-2503-02318 enable reasoning in LALMs through supervised fine-tuning (SFT) on a large-scale dataset with CoT annotations, providing the first evidence that CoT is effective for solving complex audio understanding problems. Building on this, DBLP:journals/corr/abs-2503-11197 introduces group relative policy optimization (GRPO) into LALMs for the first time, surpassing the former with less data and offering initial comparisons across different prompts. These studies focus either on SFT or GRPO, and both apply reasoning to all questions, but they lack in-depth analysis of the differences between the two approaches. Therefore, we conduct a detailed analysis to determine under what conditions SFT and GRPO are more effective. We find that GRPO performs better on harder questions, whereas on easier ones it tends to produce redundant reasoning and performs slightly worse than SFT. Furthermore, we analyze GRPO under forced reasoning (explicit prompt) and without forced reasoning (implicit prompt), and observe that the forced reasoning models maintain a clear advantage on harder questions, while the two settings perform similarly on easier ones. Taken together, these findings indicate that achieving efficient reasoning in LALMs requires adapting reasoning length to problem difficulty.
@@
Regarding reasoning length and efficiency, DBLP:journals/corr/abs-2503-21614 highlight issues such as redundancy and overthinking. RL-based studies on LLMs (DBLP:journals/corr/abs-2502-04463; DBLP:journals/corr/abs-2503-04697) design length-penalty rewards but rely on fixed thresholds that overlook question types and difficulty levels. In the audio domain, wu2025audio introduces a “when to think” mechanism that guides the model on whether reasoning is necessary. However, for the samples that still perform reasoning, it lacks a fine-grained mechanism to modulate how much to reason.Therefore, an effective and efficient LALM should reason smartly by adapting its reasoning depth to the problem’s complexity, achieving short reasoning for simple questions and deeper reasoning for difficult ones. Based on this analysis and prior work, we propose a new length-based reward function that no longer depends on fixed thresholds. In addition, we introduce two difficulty-adaptive standards to complement this reward, enabling reasoning length to align more appropriately with question difficulty. Both approaches achieve strong performance on the MMAU benchmark (DBLP:conf/iclr/SakshiTKSSNDGM25), particularly on harder questions, while also producing much shorter reasoning than direct GRPO models, thereby greatly improving reasoning efficiency. Furthermore, we conduct a qualitative case study on models from the main experiments, providing a detailed analysis of the reasoning structures in their outputs.
@@ 2 Deep Analysis of different methods for LALMs
LALMs have already demonstrated strong capability in addressing basic understanding tasks. Recent studies (DBLP:journals/corr/abs-2503-02318; DBLP:journals/corr/abs-2503-11197) focus on enhancing their ability to solve complex problems through reasoning. However, while these works adopt different implementation methods, they lack in-depth analysis of the differences between approaches. Therefore, in this section, we conduct detailed analysis from two perspectives: which is more effective under various conditions between SFT and GRPO, and whether performance gains are driven by explicit or implicit prompt.
@@
Audio Grounding AudioGrounding (DBLP:conf/icassp/XuDW021) 1,805
Sound Classification VocalSound (DBLP:conf/icassp/GongYG22) 15,531
TUT2017 (DBLP:conf/dcase/MesarosHDESVRV17) 3,744
Sound Question Answering Clotho-AQA (DBLP:conf/eusipco/LippingSDV22) 6,615
AVQA (DBLP:conf/mm/Yang0DCHJ022) 36,036
@@
For the base models, we primarily adopt Qwen2-Audio-7B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2-Audio-7B-Instruct and Qwen2.5-Omni-7B  [^2] https://huggingface.co/Qwen/Qwen2.5-Omni-7B. More detailed experimental settings, including the two prompts, are provided in Appendix  A.2. At the same time, for those experiments that share common settings, we also add an additional group average in Table  2 to facilitate comparisons across groups.
@@ 3.1 Defining Model-perspective Difficulty
$\gamma=\begin{cases}0,&C\geq 6,\\ 0.5,&3\leq C<6,\\ 1,&C<3,\end{cases}\ (G=8),\ C=\sum_{i=1}^{G}c_{i}$ (1)
@@
$a^{(n)}_{j}=\mathbf{A}^{(n)}_{T,j}\ ,\ \bar{p}_{j}=\frac{1}{N}\sum_{n=1}^{N}a^{(n)}_{j}\ ,\ H=-\sum\bar{p}_{j}\log\bar{p}_{j}\ ,(j\in\mathcal{M}).$ (2)
$\gamma^{(b)}=\frac{H^{(b)}-\min_{b^{\prime}\in\mathcal{B}}H^{(b^{\prime})}}{\max_{b^{\prime}\in\mathcal{B}}H^{(b^{\prime})}-\min_{b^{\prime}\in\mathcal{B}}H^{(b^{\prime})}}\;\in\;[0,1].$ (3)
Here, $\mathbf{A}$ denotes the complete attention matrix, $T$ represents the number of tokens, $\mathcal{M}$ indicates the indices corresponding to the audio attention part, $H$ denotes the computed entropy, and $\mathcal{B}$ is the batch size. The final $\gamma^{(b)}$ represents the difficulty value of the $b$-th sample within a batch. Compared with the first method, this approach provides a more fine-grained division of difficulty, no longer limited to three categories, but instead yielding a continuous difficulty range.
@@ 3.2 Difficulty-Adaptive Length-Based Reward
Figure 1: Curves of GRDR and GA2DR with normalized length.
After defining question difficulty, the next step is to link difficulty with reasoning length in an effective way. Our core idea is short reasoning for simple questions and long reasoning for difficult ones, with reward values that change dynamically rather than staying fixed. Thus, we design a rule-based reward using a negative exponential function, with the corresponding curve shown in Figure  1. Specifically, it smoothly adjusts reward values according to the ratio between the current reasoning length and the model’s maximum reasoning length. The decision between reward and penalty also depends on correctness. Incorrect samples always receive a penalty that decreases as reasoning length grows, encouraging further reasoning. Once a sample is answered correctly at a certain stage, the penalty switches to a positive reward, with the reward value increasing as reasoning length shortens, encouraging concise reasoning that retains only the core solution steps. For difficulty, specifically in GRDR (plot a), the difficulty levels are divided into three categories corresponding to three curves. Along the correct reward axis, difficulty increases from bottom to top, while on the wrong axis, it decreases accordingly. In contrast, GA2DR (plot b) does not fix discrete difficulty levels but allows the curve exponent to vary continuously within the region. Both follow the same trend in which simple questions correspond to steeper curves and difficult ones to flatter curves, meaning that under the same reward value, difficult questions lead to longer reasoning lengths, and under the same reasoning length, simple questions yield smaller rewards. This achieves short reasoning for simple tasks and longer reasoning for difficult ones. The detailed calculation is as follows:
@@ 4 Experiments
Table 3: The performance of models on MMAU-test-mini with different base models and reward settings, reported under both the human-perspective difficulty annotations and the assigned model-perspective difficulty annotations. Here, $\dagger$ denotes the model-perspective difficulty annotations and $\ast$ denotes a leading proprietary model.
@@
$\text{Gemini2.5-Pro-0506}^{\ast}$ [^1] https://ai.google.dev/gemini-api/docs/models#gemini-2.5-pro 70.57 65.26 62.16 52.32 95.82 77.64 57.47 55.60 12.35 66.00
@@
Table 4: Evaluations on MMAU-v0515. Here, $\ast$ denotes a leading proprietary model.
Models MMAU-Test-Mini-v0515
Sound Music Speech Avg
Baseline Models
Qwen2-Audio 62.16 62.27 55.55 60.00
Qwen2.5-Omni 74.17 65.26 61.56 67.00
Kimi-Audio 78.97 60.47 66.96 68.80
$\text{Gemini2.5-Pro}^{\ast}$ 76.57 73.95 80.78 77.10
Based On Qwen2.5-Omni-7B
GRPO 84.08 69.46 74.17 75.90
+ TR 83.78 70.65 74.47 76.30
+ GRDR 83.48 70.35 75.97 76.60
+ GA2DR 83.18 71.55 75.67 76.80
Table 5: Evaluations on MMAR. Here, $\ast$ denotes a leading proprietary model.
Models MMAR
Sound Music Speech Avg
Baseline Models
Qwen2-Audio 33.33 24.27 32.31 30.00
Qwen2.5-Omni 58.79 40.78 59.86 56.70
Kimi-Audio 57.57 45.63 63.26 59.00
$\text{Gemini2.5-Pro}^{\ast}$ 73.33 64.07 88.77 80.50
Based On Qwen2.5-Omni-7B
GRPO 60.00 48.05 62.24 59.90
+ TR 64.84 49.51 63.94 61.90
+ GRDR 61.21 51.94 65.30 61.20
+ GA2DR 64.84 54.85 65.30 62.90
Setup.
In this section, our experiments are mainly based on Qwen2-Audio-7B-Instruct and Qwen2.5-Omni-7B. All models are fine-tuned using LoRA with Prompt2. Most evaluations are conducted on MMAU-test-mini with ACC as the metric. And we also evaluate our proposed methods on MMAU-test-mini (v05.15.25), which improves the Q&A formulation and enhances the quality of the audio itself compared to the previous version, on MMAR (DBLP:journals/corr/abs-2505-13032), a benchmark designed to assess the deep reasoning capabilities of Audio Language Models (ALMs) in complex settings that span large-scale multitask, multimodal, and multilingual scenarios. Other details can be found in Appendix  A.2.
@@ 4.1 Main Results
Table  3 mainly presents three parts: the performance of four baseline models and our proposed rewards on Qwen2-Audio-7B-Instruct and Qwen2.5-Omni-7B. We further report results on MMAU-v0515 and MMAR, together with the base model on these benchmarks, shown in Table  5 and Table  5. TR denotes the basic Truncation Reward (DBLP:journals/corr/abs-2505-15612), with its formula provided in Appendix  A.2. In addition, we extend MMAU-test-mini with model-perspective difficulty labels annotated by the four baseline models in Table  3, with details given in Appendix  A.3.
@@
First, we analyze the four baseline models. Their overall performance reflects different capability levels and the typical range of most systems. However, they differ markedly on hard questions. Even Gemini2.5-Pro, which is nearly the strongest on average, performs poorly on hard items, while Qwen2-Audio, the weakest on average, shows a clear advantage. We attribute this to the combined effects of the LLM backbone strength and the number of supported modalities. For easy questions, the LLM backbone is dominant because these items require minimal audio understanding and rely mainly on text comprehension. In contrast, hard questions demand stronger audio interpretation, where pure LALMs often outperform Omni-style models that integrate more modalities. This pattern is evident in Qwen2.5-Omni and Kimi-Audio, which share a “Whisper + Qwen2.5” architecture but still exhibit a noticeable gap on hard questions.
Second, we compare our proposed GRDR and GA2DR with TR. For Qwen2-Audio-7B-Instruct, the three methods achieve similar overall performance, with GRDR performing best and TR worst, confirming the effectiveness of our approach. TR shows slightly better performance on medium questions, but GRDR is clearly stronger on hard ones. For Qwen2.5-Omni-7B, both of our methods deliver clear overall gains. They substantially outperform TR on medium and hard questions, while maintaining comparable results on easy ones. This shows that our methods effectively utilize question difficulty to assign appropriate rewards, promote deeper reasoning on hard questions, and improve performance on challenging tasks. Unlike TR, which uses a fixed truncation length and only constrains part of the samples, our reward designs offer more balanced treatment across difficulty levels, leading to stronger overall results.
Third, we compare GRDR and GA2DR, focusing on the results based on Qwen2.5-Omni-7B. The two methods differ in that GRDR is outcome-oriented, whereas GA2DR is process-oriented. GRDR performs better on easy questions, GA2DR is stronger on medium ones, and their performance is similar on hard questions. We believe this is due to their different difficulty definitions: GRDR uses only three levels, while GA2DR applies a finer, unconstrained difficulty scale, which particularly benefits medium questions. For easy and hard questions, GA2DR is slightly weaker because normalization may place samples with similar audio attention entropy into different difficulty bins, diminishing its advantage.
@@
Performance across Benchmarks.
To further evaluate our model in broader and more reasoning-intensive scenarios, we conduct additional evaluations on MMAU-v0515 and MMAR, which also help verify the generalization ability of our methods. In terms of overall performance, our two methods still achieve the best results on MMAU-v0515 compared with all baselines. On MMAR, however, the two methods behave differently: GA2DR maintains a clear performance lead, whereas GRDR falls behind the TR method.
This difference can be explained by the nature of the two approaches. GRDR is outcome-oriented, while GA2DR is process-oriented. GRDR is highly susceptible to noise in rollout samples, making it prone to reward hacking. As a result, it performs reasonably well on the relatively simpler MMAU benchmarks but degrades significantly on the more challenging MMAR benchmark. In contrast, GA2DR is unaffected by such noise because it determines difficulty solely based on the model’s current behavior and the audio characteristics of each question. This removes much of the randomness and reduces the likelihood of reward hacking, leading to more stable performance across benchmarks and consistently strong results on more challenging tasks. We also conduct additional experiments to further verify generalization. The results confirm that GRDR is indeed affected by random noise introduced during rollout; however, once this issue is mitigated through appropriate constraints, its performance still surpasses all baseline methods. Detailed results and analysis are provided in Appendix  A.4.
@@ 4.2 Analysis of Reasoning Length across Models
In this subsection, we mainly compare reasoning length from both difficulty perspectives, as shown in Figure  2, focusing on the four baseline models and all models based on Qwen2.5-Omni-7B. Because the length gaps between models are relatively large, we present log-scaled lengths in this part. We also provide Figure  3, which shows three Qwen2.5-Omni-7B–based models—direct GRPO and our two proposed rewards. To better illustrate how our methods adjust reasoning length across different difficulty levels, this figure uses the actual token counts without log-scaling. In addition, detailed length statistics are provided in Appendix  A.5.
@@
Figure 3: The trend of average reasoning length for direct GRPO and our two proposed methods on MMAU-Test-Mini, evaluated under both human-perspective and model-perspective difficulty. Length is measured directly in tokens without any logarithmic transformation.
@@
4.3 Ablation study: Results of single length-based reward and $k$ settings
In this section, we conduct an in-depth ablation study on our proposed difficulty-adaptive length-based reward. Specifically, we remove the mechanism in Equation  5 that determines different reward curves based on question difficulty, and replace $k(\lambda)$ in Equation 4 with a fixed value $k$, meaning that all questions share the same reward curve. We also experiment with different settings of $k$ to further validate the effectiveness of our approach. The main evaluation is performed on MMAU-Test-Mini in Table  6, and additional results on MMAU-v0515 and MMAR are also provided in Table 8 and Table  8.
Table 6: The performance of models trained with the single length-based reward setting. Here, $\dagger$ denotes the model-perspective difficulty annotations.
Models MMAU-Test-Mini
GRPO Based On Qwen2.5-Omni-7B (With Length-based Reward Only)
k=2 73.27 67.06 67.86 59.69 94.49 77.84 58.41 61.63 27.41 69.40
k=6 72.37 66.46 68.76 56.20 93.92 79.21 57.00 61.63 28.95 69.20
k=10 72.37 70.06 66.36 58.91 93.92 79.60 56.07 59.48 31.27 69.60
Table 7: Evaluations on MMAU-v0515.
Models MMAU-Test-Mini-v0515
Sound Music Speech Avg
GRPO Based On Qwen2.5-Omni-7B
k=2 81.38 73.05 74.77 76.40
k=6 82.88 71.55 74.77 76.40
k=10 81.98 71.25 76.57 76.60
Table 8: Evaluations on MMAR.
Models MMAR
Sound Music Speech Avg
GRPO Based On Qwen2.5-Omni-7B
k=2 61.81 50.97 63.26 61.40
k=6 60.60 52.91 64.96 61.80
k=10 61.81 52.91 64.62 62.20
From the perspective of overall performance, all fixed k-value settings outperform direct GRPO and the traditional TR method. This demonstrates the effectiveness of our length-based reward and shows that these samples indeed require length optimization—direct GRPO tends to produce either redundant or insufficient reasoning. It also indicates that, compared with TR, our reward design can dynamically adjust rewards based on length rather than relying on a fixed threshold, resulting in better optimization. Comparing with the difficulty-adaptive methods GRDR and GA2DR, most fixed k settings perform worse, further confirming the validity and effectiveness of incorporating adaptive difficulty. Across the two MMAU benchmarks, performance fluctuates noticeably across all three difficulty levels under different k values, showing that applying a single reward curve to all questions leads to imbalance and reduced overall performance. Overall, these experiments further validate the effectiveness of our difficulty-adaptive length-based reward.